# Phase Transition of Spectral Fluctuations in Large Gram Matrices with a Variance Profile: A Unified Framework for Sparse CLTs

### v1

@@ 5.2 Outage probability analysis
The MIMO channel is generated according to (6), where the diagonal entries are drawn as $d_{i}\sim\mathrm{Uniform}(1,2)$ for $i=1,\ldots,N_{r}$ and $\tilde{d}_{j}\sim\mathrm{Uniform}(1,2)$ for $j=1,\ldots,N_{t}$. The random matrix $\mathbf{X}$ consists of i.i.d. entries $x_{ij}\sim\mathcal{CN}(0,1)$. Also, two regimes of $q$ are examined, corresponding to $q=N_{t}^{1/3}$ (high-sparsity regime) and $q=0.5\,N_{t}^{1/2}$ (moderate-sparsity regime). Unless otherwise specified, we set $(N_{r},N_{t})=(512,1024)$ and average the results over $2,000$ independent channel realizations.
Figure 3 compares the empirical and theoretical distributions of the normalized statistic $T_{\log}$ for $\sigma=1$ under two sparsity regimes. In the moderate-sparsity case of $q=0.5N_{t}^{1/2}$, the histogram of $T_{\log}$ nearly coincides with the standard normal density. A similar observation holds for the high-sparsity regime of $q=N_{t}^{1/3}$. Figure 4 shows the empirical and theoretical outage probabilities versus the transmission rate $R$ for different signal-to-noise ratios (SNRs). The SNR (in dB) is related to the noise variance by $\mathrm{SNR}_{\mathrm{dB}}=10\log_{10}(1/\sigma^{2})$, i.e., $\sigma^{2}=10^{-\mathrm{SNR}_{\mathrm{dB}}/10}$. Three SNR values are considered in the simulations, namely 0, $0.2$, and $0.4~\mathrm{dB}$. The analytical curves closely match the Monte Carlo results in both sparsity regimes, and a clear rightward shift is observed with increasing SNR, indicating improved achievable rates.
@@ Appendix A Proofs
Proof of (2). We consider $x_{ij}=b_{ij}w_{ij}/\sqrt{s}$, where $b_{ij}\sim\operatorname{Bernoulli}(1,s)$ and $w_{ij}$ are centered with $\mathbb{E}|w_{ij}|^{2}=\sigma_{ij}^{2}$ and finite $4+\epsilon$ moments. Unlike the dense model in [10], sparsity introduces a significant challenge: in the high-sparsity regime ($s\to 0$), the $4+\epsilon$ moments of $x_{ij}$ may diverge, and classical resolvent estimates for quadratic forms no longer apply directly. When $s$ is fixed and bounded away from zero (moderate sparsity), classical techniques can be adapted. Our main contribution is the extension of these results to the high-sparsity regime, which relies on refined resolvent-based bounds for quadratic forms under diverging entrywise moments.
@@
$\left\|\frac{1}{\tilde{t}_{j}(z)}-\frac{1}{\tilde{p}_{j}(z)}\right\|_{2+\epsilon/2}\leq\frac{K_{1}|z|\sigma_{\max}^{2}}{|\Im z|^{2}}\sup_{k}\left\|\frac{1}{t_{k}(z)}-\frac{1}{p_{k}(z)}\right\|_{2+\epsilon/2}+\frac{K_{2}|z|}{q}.$ (12)
@@
Moreover, if $q\to\infty$ sufficiently fast so that $\sum_{n}q^{-2-2+\epsilon/2}<\infty$, then Borel–Cantelli yields almost sure convergence along subsequences. Finally, by analyticity and the fact that the resolvent traces form a normal family on any compact subset of $\mathbb{C}\setminus\mathbb{R}$ (see [10, Section 6.3]), the convergence extends uniformly on such sets.
@@
To characterize the asymptotic distribution in Theorem 3.3, we first introduce an auxiliary result that provides the key quantities $U_{j}$ and $\tilde{U}_{i}$. The proof is provided in the supplementary material.
@@
Consider a variance profile $(\sigma_{ij}^{2})$ satisfying Assumptions 2.2, and let $t_{i}$, $\tilde{t}_{j}$ be as defined in Theorem 3.1. Fix $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
@@
    For $1\leq j\leq n$, the quantity $U_{j}$ admits the representation
    $U_{j}=\frac{1}{z^{2}\tilde{t}_{j}^{2}(z)}\breve{y}_{j,j},$
    where $\breve{y}_{j,j}$ is determined by the $n$-dimensional linear system
    $\breve{y}_{l,j}=\sum_{i\neq j}^{n}a_{li}\breve{y}_{i,j}+na_{lj},\quad 1\leq l\leq n,$ (13)
    in the unknowns $\{\breve{y}_{l,j}:1\leq l,j\leq n\}$. This system has a unique solution for all sufficiently large $n$.
@@
    Similarly, for $1\leq i\leq p$, the quantity $\tilde{U}_{i}$ satisfies
    $\tilde{U}_{i}=\frac{1}{z^{2}t_{i}^{2}(z)}\tilde{y}_{i,i},$
    where $\tilde{y}_{i,i}$ is obtained from the $p$-dimensional linear system
    $\tilde{y}_{i,m}=\sum_{l\neq m}^{p}\acute{a}_{il}\tilde{y}_{l,m}+n\acute{a}_{im},\quad 1\leq i,m\leq p,$ (14)
@@
    $\acute{a}_{im}=\frac{1}{n^{2}}\frac{\operatorname{Tr}\left(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}(z)\widetilde{\mathbf{\Sigma}}_{m}\widetilde{\mathbf{T}}(z)\right)}{\left(1+\frac{1}{n}\operatorname{Tr}\left(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}(z)\right)\right)^{2}}.$
@@
Let $\widehat{L}^{c}_{n}(f)$ and $\widetilde{L}^{c}_{n}(f)$ denote the analogues of $L^{c}_{n}(f)$ obtained by replacing $\mathbf{S}_{n}$ with $\widehat{\mathbf{S}}_{n}$ and $\widetilde{\mathbf{S}}_{n}$, respectively. Let $\lambda_{i}^{\mathbf{A}}$ be the $i$th largest eigenvalue of a Hermitian matrix $\mathbf{A}$. Using the argument in [4, Corollary A.42] and Cauchy–Schwarz’s inequality, we obtain
@@
where $\tilde{\nu}_{8+\epsilon}$ denotes a finite constant depending on the $8+\epsilon$ moment of the underlying distribution. Moreover, $\delta_{n}$ satisfies $\lim_{n\rightarrow\infty}1/(\delta_{n}q)^{4}\sum_{i,j}\mathrm{E}\left|w_{ij}\right|^{4}I\left(\left|w_{ij}\right|\geq\delta_{n}q\right)=0$, $\delta_{n}\downarrow 0$, and $\delta_{n}q\uparrow\infty$. All subsequent analysis is carried out under these bounded-moment conditions.
@@
$\int f(x)\mathrm{d}Q(x)=-\frac{1}{2\pi i}\oint f(z)m_{Q}(z)\mathrm{d}z$

### v2

@@ 5.2 Outage probability analysis
The MIMO channel is generated according to (7), where the diagonal entries are drawn as $d_{i}\sim\text{Uniform}(1,2)$ for $i=1,\ldots,N_{r}$ and $\tilde{d}_{j}\sim\text{Uniform}(1,2)$ for $j=1,\ldots,N_{t}$. The random matrix $\mathbf{X}$ consists of i.i.d. entries $x_{ij}\sim\mathcal{CN}(0,1)$. Also, two cases of $q$ are considered, corresponding to $q=0.5N_{t}^{1/2}$ (moderate-sparsity regime) and $q=N_{t}^{1/3}$ (high-sparsity regime). Unless otherwise specified, we set $(N_{r},N_{t})=(512,1024)$ and average the results over $2,000$ independent channel realizations.
Figure 3 compares the empirical and theoretical distributions of the normalized statistic $T_{\log}$ for $\sigma=1$ under two sparsity regimes. In the moderate-sparsity case of $q=0.5N_{t}^{1/2}$, the histogram of $T_{\log}$ nearly coincides with the standard normal density. A similar observation holds for the high-sparsity regime of $q=N_{t}^{1/3}$. Figure 4 shows the empirical and theoretical outage probabilities versus the transmission rate $R$ for different signal-to-noise ratios (SNRs). The SNR (in dB) is related to the noise variance by $\text{SNR}_{\text{dB}}=10\log_{10}(1/\sigma^{2})$, i.e., $\sigma^{2}=10^{-\text{SNR}_{\text{dB}}/10}$. Three SNR values are considered in the simulations, namely 0, $0.2$, and $0.4~\text{dB}$. The analytical curves closely match the Monte Carlo results in both sparsity regimes, and a clear rightward shift is observed with increasing SNR, indicating improved achievable rates.
@@ Appendix A Proofs
Proof of (2). We consider $x_{ij}=b_{ij}w_{ij}/\sqrt{s}$, where $b_{ij}\sim\operatorname{Bernoulli}(1,s)$ and $w_{ij}$ are centered with $\mathbb{E}|w_{ij}|^{2}=\sigma_{ij}^{2}$ and finite $4+\epsilon$ moments. Unlike the dense model in [10], sparsity introduces a significant challenge: in the high-sparsity regime ($s\to 0$), the $4+\epsilon$ moments of $x_{ij}$ may diverge, and classical resolvent estimates for quadratic forms no longer apply directly. When $s$ is fixed and bounded away from zero (moderate sparsity), classical techniques can be adapted. Our main contribution is the extension of these results to the high-sparsity regime, which relies on refined resolvent-based bounds for quadratic forms under diverging entry-wise moments.
@@
$\left\|\frac{1}{\tilde{t}_{j}(z)}-\frac{1}{\tilde{p}_{j}(z)}\right\|_{2+\epsilon/2}\leq\frac{K_{1}^{\prime}|z|\sigma_{\max}^{2}}{|\Im z|^{2}}\sup_{k}\left\|\frac{1}{t_{k}(z)}-\frac{1}{p_{k}(z)}\right\|_{2+\epsilon/2}+\frac{K_{2}^{\prime}|z|}{q}.$ (13)
@@
Moreover, if $q\to\infty$ sufficiently fast so that $\sum_{n}q^{-2-2+\epsilon/2}<\infty$, then Borel–Cantelli yields almost sure convergence along subsequences. Finally, by analyticity and the fact that the resolvent traces form a normal family on any compact subset of $\mathbb{C}\setminus\mathbb{R}^{+}$ (see [10, Section 6.3]), the convergence extends uniformly on such sets.
@@
To characterize the asymptotic distribution in Theorem 3.3, we first introduce an auxiliary lemma that provides the key quantities $U_{j}(z)$ and $\tilde{U}_{i}(z)$. The proof is provided in the supplementary material.
@@
Consider a variance profile $(\sigma_{ij}^{2})$ satisfying Assumptions 2.2-2.3 , and let $t_{i}$, $\tilde{t}_{j}$ be as defined in Theorem 3.1. Fix $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
@@
    For $1\leq j\leq n$, the quantity $U_{j}(z)$ admits the representation
    $U_{j}(z)=\frac{1}{z^{2}\tilde{t}_{j}^{2}(z)}\breve{y}_{j,j}(z),$
    where $\breve{y}_{j,j}(z)$ is determined by the $n$-dimensional linear system
    $\breve{y}_{l,j}(z)=\sum_{i\neq j}^{n}a_{li}(z,z)\breve{y}_{i,j}(z)+na_{lj}(z,z),\quad 1\leq l,j\leq n,$ (14)
    in the unknowns $\{\breve{y}_{l,j}(z):1\leq l,j\leq n\}$. This system has a unique solution for all sufficiently large $n$.
@@
    Similarly, for $1\leq i\leq p$, the quantity $\tilde{U}_{i}(z)$ satisfies
    $\tilde{U}_{i}(z)=\frac{1}{z^{2}t_{i}^{2}(z)}\tilde{y}_{i,i}(z),$
    where $\tilde{y}_{i,i}(z)$ is obtained from the $p$-dimensional linear system
    $\tilde{y}_{i,m}(z)=\sum_{l\neq m}^{p}\acute{a}_{il}(z)\tilde{y}_{l,m}(z)+n\acute{a}_{im}(z),\quad 1\leq i,m\leq p,$ (15)
@@
    $\acute{a}_{im}(z)=\frac{1}{n^{2}}\frac{\operatorname{Tr}\left(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}(z)\widetilde{\mathbf{\Sigma}}_{m}\widetilde{\mathbf{T}}(z)\right)}{\left(1+\frac{1}{n}\operatorname{Tr}\left(\widetilde{\mathbf{\Sigma}}_{i}\widetilde{\mathbf{T}}(z)\right)\right)^{2}}.$
@@
Let $\widehat{L}^{c}_{n}(f)$ and $\widetilde{L}^{c}_{n}(f)$ denote the analogues of $L^{c}_{n}(f)$ obtained by replacing $\mathbf{S}_{n}$ with $\widehat{\mathbf{S}}_{n}$ and $\widetilde{\mathbf{S}}_{n}$, respectively. Let $\lambda_{i}^{\mathbf{A}}$ be the $i$-th largest eigenvalue of a Hermitian matrix $\mathbf{A}$. Using the argument in [2, Corollary A.42] and Cauchy–Schwarz’s inequality, we obtain
@@
where $\tilde{\nu}_{8+\epsilon}$ denotes a finite constant depending on the $8+\epsilon$ moment of the underlying distribution. Moreover, $\delta_{n}$ satisfies $\lim_{n\rightarrow\infty}1/(\delta_{n}q)^{4}\sum_{i,j}\mathbb{E}\left|w_{ij}\right|^{4}I\left(\left|w_{ij}\right|\geq\delta_{n}q\right)=0$, $\delta_{n}\downarrow 0$, and $\delta_{n}q\uparrow\infty$. All subsequent analysis is carried out under these bounded-moment conditions.
@@
$\int f(x)\mathrm{d}Q(x)=-\frac{1}{2\pi i}\oint f(z)m_{Q}(z)\mathrm{d}z,$